OpenStudy (anonymous):

How to do an equation like this 3/6/2

OpenStudy (anonymous):

\[\frac{x}{\frac{m}{n}} = \frac{x \times n}{m}\] Getting this formula or not??

mathslover (mathslover):

can u write :3/6/2 = 3/6/(2/1) = 3/6*2/1 can u do it now ?

OpenStudy (anonymous):

Yea thanks you